what is 4(2a-2)=-2(1-5a)
Answer by mmm4444bot(95)

Hello There:
4*(2*a - 2) = -2*(1 - 5*a)
(The asterisks mean multiplication.)
Start by using the distributive property to get rid of the parentheses. In other words, both the 2*a and the -2 get multiplied by 4. On the right side, both the 1 and the -5*a get multiplied by -2.
8*a - 8 = -2 + 10*a
Next, we want to get the two terms that contain the variable 'a' together on one side of the equals sign, so that we can combine them into one term. We can either subtract 8*a from both sides or subtract 10*a from both sides.
I'll choose to subtract 8*a from both sides.
-8 = -2 + 2*a
Next, we combine the constants. Add 2 to both sides.
-6 = 2*a
Finally, we solve for a by dividing both sides by 2 to isolate the 'a'.
a = -3
You can check this result by substituting -3 for 'a' in the original equation, do the arithmetic on each side of the equals sign, and see if you come up with a true statement.
~ Mark